OpenStudy (anonymous):

Find the quotient. Write your answer in standard form. (2+i) / (2-i) ***1-i?***

OpenStudy (cwrw238):

Multiply top and bottom of the fraction by the conjugate of 2 - i

OpenStudy (anonymous):

That would give me \[\frac{ 3+4i }{5 }\] right?

OpenStudy (anonymous):

I see now... Haha thank you guys! The answer would be \[\frac{ 3 }{ 5} + \frac{ 4 }{ 5 }i\]Right?

zepdrix (zepdrix):

yay good job \c:/